Susan Bertrees can invest $12,000 at 2% interest compounded twice a year or compounded quarterly. If either investment is for 5 years, which investment results in more interest? How much more interest is yielded by the better investment?


1
Expert's answer
2021-10-18T14:36:02-0400

compounded twice a year:

"FV_1=12000(1+0.02/2)^{2\cdot5}=\$13255.47"


compounded quarterly:

"FV_2=12000(1+0.02/4)^{4\cdot5}=\$13258.75"


"FV_2-FV_1=13258.75-13255.47=\$3.28"